OpenStudy (anonymous):

solve for t 16(t-1) = t(t+8) please show your work, so I can work through it.

OpenStudy (anonymous):

When I try to solve it I end up with a negative root (complex solution)

OpenStudy (anonymous):

distribute them 16t - 16 = t^2 + 8t t^2 - 8t + 16 = 0 (t-4)^2 = 0 t = 4

OpenStudy (anonymous):

ohh (Facepalm) thank you

OpenStudy (anonymous):

i could've just moved the -16 over...